# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ III Quantization of the regularized quasinormal modes
(×$-(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg{]}\ .$
@@
(×$\cdot(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg{]}\ ,$
showing that it is not altered by the presence of the PMLs and reflects the correlation function of the real system without PMLs. Our modal theory thus reproduces the correct commutators and vacuum correlation functions of the exact theory. The magnetic field operator can be expanded using the same bosonic modal operators:
@@ 2. Combination of gain/loss operators
just as we have seen when we included the gain components of the PMLs in the loss operators of Eqs. (III) and (III). However, caution is required at this point, because the vacuum fluctuations and the corresponding pumping of the modes due to the gain components of the resonator have a physical influence that generally cannot be neglected, and which becomes clear in the effective master equation derived in § III. As discussed in Ref. Franke, Ren, and Hughes (2022), two conditions must be met for the gain and loss operators to be combined. Firstly, the combined matrices $P^{\mathrm{\lambda}}_{nn^{\prime}}$ must be positive-definite, which is required for the symmetrization of the modal operators. Secondly, as already pointed out when treating the gain components of the PMLs, the combined operators $\hat{\xi}_{n,\lambda}$ do not have the same vacuum state |$\ket{0}$ as the continuous variables $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$. Applying the modal number operators (without the contributions from the PML gain components) to the vacuum state leads to $\sum_{n}\hat{a}^{\dagger}_{n,\lambda}\hat{a}_{n,\lambda}\ket{0}=\sum_{n}n^{0}_{n,\lambda}\ket{0}$ with $\sum_{n}n^{0}_{n,\lambda}=\mathrm{Tr}\{P^{\lambda,\mathrm{G}*}(P^{\lambda})^{-1}\}$. Consequently, the vacuum state of the complete system can only be used approximately as a vacuum state of the modal Fock space if $\sum_{n}n^{0}_{n,\lambda}\ll 1$, which holds whenever the losses in the system dominate the gain (i.e., $P^{\lambda,\mathrm{L}}_{nn}\gg P^{\lambda,\mathrm{G}}_{nn}$) and the off-diagonal elements $P^{\lambda,\mathrm{G}}_{nn^{\prime}}$ are small. In this case, the expanded electric field operator would read
@@
in which the symmetrized modes are as defined in Eq. (48) with the replacement $P^{\lambda,\mathrm{L}}\rightarrow P^{\lambda}$. Note that in this approximation, quantum effects due to the vacuum fluctuations caused by the presence of the gain regions are neglected, however, they still influence the dynamics as their presence alters the modes and the Green’s function of the system.
@@ 3. Combination of all four operators
It is also worth mentioning that in our approach, the limit of lossless/gainless materials, (i.e., $\zeta_{\mathrm{i}}\rightarrow 0$), can be taken from the beginning, since we map the problem onto a compact domain $\Omega\subset\mathbb{R}^{3}$. This means that in this case, we can interchange the limit and the integration in the symmetrization matrix elements $P^{\lambda,\mathrm{L}}_{nn^{\prime}}$, and consequently only the PML region $\Omega_{\mathrm{PML}}$ contributes, which corresponds to the radiation loss in Ref. Franke et al. (2020b).
@@ IV Effective Lindblad master equation
In this section, we derive an effective Lindblad master equation for the electromagnetic modes and the additional coupling to a QE, described as a two-level system (TLS). We will start by considering the dynamics of the modes only in the most general scenario, in which they are characterized by the operators $\hat{a}_{n,\lambda,X}$. In the Heisenberg picture, the dynamics of each modal operator is determined by the Heisenberg equation of motion $\mathrm{d}\hat{O}/\mathrm{d}t=\mathrm{i}[\hat{H},\hat{O}]/\hbar$, leading to Franke, Ren, and Hughes (2022)
$\frac{\mathrm{d}\hat{a}_{n,\lambda,\mathrm{L}}}{\mathrm{d}t}=\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\hat{a}_{n,\lambda,\mathrm{L}}\right]-(\tilde{\chi}^{\lambda,\mathrm{L},-}_{nn^{\prime}}-\tilde{\chi}^{\lambda,\mathrm{PML},-}_{nn^{\prime}})\hat{a}_{n^{\prime},\lambda,\mathrm{L}}$
$+\hat{F}_{n,\lambda,\mathrm{L}}(t)+\hat{F}^{\dagger}_{n,\lambda,\mathrm{PML}}(t)\ ,$ (63)
@@
$\tilde{\chi}^{\lambda,\mathrm{L},-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}(P^{\lambda,\mathrm{L}})^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})\tilde{P}^{\lambda,\mathrm{L}}_{n^{\prime\prime}n^{\prime\prime\prime}}(P^{\lambda,\mathrm{L}})^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ ,$ (67)
$\tilde{\chi}^{\lambda,\mathrm{PML},-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}(P^{\lambda,\mathrm{L}})^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})\tilde{P}^{\lambda,\mathrm{PML}}_{n^{\prime\prime}n^{\prime\prime\prime}}(P^{\lambda,\mathrm{L}})^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ ,$ (68)
$\chi^{\lambda,\mathrm{G},-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}(P^{\lambda,\mathrm{G}})^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})P^{\lambda,\mathrm{G}}_{n^{\prime\prime}n^{\prime\prime\prime}}(P^{\lambda,\mathrm{G}})^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ .$ (69)
@@
The $\hat{F}_{n,\lambda,\mathrm{L/G/PML}}(t)$ are additional operator terms that still depend on the continuous variables $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$, and which cannot be expressed simply in terms of the modal operators. From the classical QNM theory, it is clear that the amplitudes of the freely evolving modes are oscillating with the complex frequencies $\tilde{\omega}_{n}$, as can be seen in Eq. (20), meaning the expectation values of the modal operators should satisfy $\langle\hat{\xi}_{n,\lambda,\mathrm{L}}(t)\rangle=\langle\hat{\xi}_{\mu,\lambda,\mathrm{L}}(0)\rangle\mathrm{e}^{-\mathrm{i}\tilde{\omega}_{n}t}$ and $\langle\hat{\xi}_{n,\lambda,\mathrm{G/PML}}(t)\rangle=\langle\hat{\xi}_{n,\lambda,\mathrm{G/PML}}(0)\rangle\mathrm{e}^{+\mathrm{i}\tilde{\omega}^{*}_{n}t}$. From that requirement, it is clear that $\langle\hat{F}^{(\dagger)}_{n,\lambda,\mathrm{L/G/PML}}(t)\rangle=0$. Furthermore, the peaked structure was factored out in the spectral representations of the operators $\hat{F}_{n,\lambda,\mathrm{L/G/PML}}(t)$ due to terms of the form $(\tilde{\omega}_{\mu}-\omega)A_{\mu}(\omega)=\varepsilon_{0}c^{2}/\tilde{\omega}_{\mu}$. We thus make a Markov approximation and treat them as white quantum noise forces induced by the dissipation/gain and that ensure the preservation of the commutation relations. This procedure is similar to the reaction coordinate mapping, in which the coupling of a TLS to a structured bath having a strongly peaked spectral density is mapped to an interaction with a discrete mode and a residual continuum that can then be treated within the Markov approximation Binder et al. (2018); Anto-Sztrikacs and Segal (2021). Note that in the equation for the modal loss operators $\hat{a}_{n,\lambda,\mathrm{L}}$ two noise forces $\hat{F}^{(\dagger)}_{m,\lambda,\mathrm{L/PML}}$ appear that counteract the loss and gain terms described by $\tilde{\chi}_{nn^{\prime}}^{\lambda,\mathrm{L/PML},-}$, caused by the overall loss in the system and the gain components of the PMLs, which we have included in the loss operators [see Eq. (III)]. In the equation for the modal gain operators $\hat{a}_{n,\lambda,\mathrm{G}}$, only one noise force $\hat{F}_{m,\lambda,\mathrm{G}}$ appears that counteracts a loss term proportional to $\chi^{\lambda,\mathrm{G}.-}_{nn^{\prime}}$. The pumping of the modes due to the gain components of the resonator are described by negative-energy contributions in the effective Hamiltonian of Eq. (65), associated to the separate modal gain operators.
@@
By invoking a quantum analogue of Itô-Stratonovich calculus, following the same procedure as in Refs.Gardiner and Collett (1985); Franke et al. (2019); Franke, Ren, and Hughes (2022), an effective Markovian master equation, corresponding to the Heisenberg-Langevin equations in Eqs. (IV) and (64), can be obtained in the form
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\rho\right]+\tilde{\chi}^{\lambda,\mathrm{L},-}_{nn^{\prime}}\mathcal{D}^{\lambda,\mathrm{L}}_{nn^{\prime}}\rho+\chi^{\lambda,\mathrm{G},-}_{nn^{\prime}}\mathcal{D}^{\lambda,\mathrm{G}}_{nn^{\prime}}\rho$ (70)
$+\tilde{\chi}^{\lambda,\mathrm{PML},-}_{nn^{\prime}}\mathcal{D}^{\lambda,\mathrm{PML}}_{nn^{\prime}}\rho\ ,$
@@
$\mathcal{D}^{\lambda,X}_{nn^{\prime}}\rho$ $=2\hat{a}_{n^{\prime},\lambda,X}\rho\hat{a}_{n,\lambda,X}^{\dagger}-\{\rho,\hat{a}_{n,\lambda,X}^{\dagger}\hat{a}_{n^{\prime},\lambda,X}\}\ ,$ (71)
$\mathcal{D}^{\lambda,\mathrm{PML}}_{nn^{\prime}}\rho$ $=2\hat{a}^{\dagger}_{n,\lambda,\mathrm{L}}\rho\hat{a}_{n^{\prime},\lambda,\mathrm{L}}-\{\rho,\hat{a}_{n^{\prime},\lambda,\mathrm{L}}\hat{a}_{n,\lambda,\mathrm{L}}^{\dagger}\}\ ,$ (72)
where the terms $\mathcal{D}^{\lambda,X}_{nn^{\prime}}\rho$ describe dissipation and $\mathcal{D}^{\lambda,\mathrm{PML}}_{nn^{\prime}}\rho$ the incoherent pumping by the gain components of the PMLs.
This is the effective master equation that governs the quantum dynamics of the modes for the most general case, featuring decomposed electric/magnetic modal operators. Note that the loss/gain present in the system necessitates a symmetrization of the modes and consequently induces a coupling between the bosonic modes in both the effective Hamiltonian and the Lindblad dissipators, (i.e., coherent and dissipative coupling Wang and Hu (2020)). The cross terms in the Lindblad dissipators capture interference effects between different decay channels, as the coupled modes dissipate into the same bath. Such cross terms were also derived for a Jaynes-Cummings model where both the QE and the cavity mode dissipate into the same bath, leading to coupled dissipators between the QE and mode operators to describe Fano-like effects Yamaguchi, Lyasota, and Yuge (2021).
To describe the coupling of the modes to a fixed QE at $\bm{r}_{\mathrm{QE}}\in\Omega\setminus\Omega_{\mathrm{PML}}$, modeled as a TLS with eigenstates |$\ket{\mathrm{g}}$, |$\ket{\mathrm{e}}$ and electric or magnetic transition dipole moment $\bm{d}/\bm{m}=\langle\mathrm{g}|\hat{\bm{d}}/\hat{\bm{m}}|\mathrm{e}\rangle$, the mQED Hamiltonian $\hat{H}_{\mathrm{EM}}$ in Eq. (29) would have to be supplemented by additional terms,
@@
in which $\sigma^{\pm}=\ket{\mathrm{e/g}}\bra{\mathrm{g/e}}$ | are the usual TLS raising and lowering operators. Note that whilst we treat both possibilities in Eq. (73) for completeness, the distinct selection rules pertaining to electric and magnetic dipole transitions will ensure that only one of $\bm{d}$ or $\bm{m}$ is non-zero in a given physical application. By using the modal expansion of the electric field operator in Eq. (50) and following the same steps as before, the master equation thereby obtained possesses the same form as in Eq. (70), but with an effective Hamiltonian
@@
and we have applied the rotating-wave approximation. It can be appreciated that the gain modal creation operators $\hat{a}^{\dagger}_{n,\lambda,\mathrm{G}}$ couple to $\sigma^{+}$ in the rotating-wave approximation since they appear in the positive-frequency parts of the fields in Eqs. (50) and (54). If the QE is subject to additional non-radiative decay channels or dephasing processes, the corresponding dissipators should be added to the master equation in Eq. (70). Fig. 2(a) offers a schematic representation of the quantum description for an exemplary resonator-QE system, based on the master equation in this general setting, where the action of the operators and all the different couplings and decay channels are labeled. Note that even though the master equation does not contain coupled dissipators between the QE and mode operators, as derived in Ref. Yamaguchi, Lyasota, and Yuge (2021), it is in principle capable of describing coherent Fano interference effects occurring in the QE decay, as the coupling of the QE with the complete set of modes captures its interaction with the electromagnetic environment completely. Such cross terms would appear if a subset of the modes were traced out, as both the QE and the remaining modes couple to these modes.
If the electric and magnetic modal operators are combined, the derivation of the master equation would proceed in a similar manner as before, with the result having the same form as Eq. (70), but with the replacements $\chi^{\lambda,\mathrm{G},\pm}\rightarrow\chi^{\mathrm{G},\pm}$, $\tilde{\chi}^{\lambda,\mathrm{L},\pm}\rightarrow\chi^{\mathrm{L},\pm}$, $\tilde{\chi}^{\lambda,\mathrm{PML},\pm}\rightarrow\chi^{\mathrm{PML},\pm}$, which are exactly defined as in Eqs. (66)-(69), but with the replacements $P^{\lambda,X}\rightarrow P^{X}$, $\tilde{P}^{\lambda,\mathrm{L/PML}}\rightarrow\tilde{P}^{\mathrm{L/PML}}$.
@@
$\frac{\mathrm{d}\hat{a}_{n}}{\mathrm{d}t}=\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\hat{a}_{n}\right]-(\tilde{\chi}_{nn^{\prime}}^{\mathrm{L},-}-\tilde{\chi}^{\mathrm{PML+G},-}_{nn^{\prime}})\hat{a}_{n^{\prime}}$ (76)
$+\hat{F}_{n,\mathrm{L}}(t)+\hat{F}^{\dagger}_{n,\mathrm{PML+G}}(t)\ ,$
@@
$\tilde{\chi}^{\mathrm{L},-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}P^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})\tilde{P}^{\mathrm{L}}_{n^{\prime\prime}n^{\prime\prime\prime}}P^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ ,$ (79)
$\tilde{\chi}^{\mathrm{PML+G},-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}P^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})$
$\hskip 54.06006pt\cdot(\tilde{P}^{\mathrm{PML}}_{n^{\prime\prime}n^{\prime\prime\prime}}+P^{\mathrm{G}}_{n^{\prime\prime}n^{\prime\prime\prime}})P^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ .$ (80)
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{i}}{\hbar}\left[H_{\mathrm{eff}},\hat{a}_{n}\right]+\tilde{\chi}_{nn^{\prime}}^{\mathrm{L},-}\mathcal{D}^{\mathrm{L}}_{nn^{\prime}}\rho+\tilde{\chi}_{nn^{\prime}}^{\mathrm{PML+G},-}\mathcal{D}^{\mathrm{PML+G}}_{nn^{\prime}}\rho\ ,$ (81)
in which the dissipators $\mathcal{D}^{\mathrm{L}}_{nn^{\prime}}$, $\mathcal{D}^{\mathrm{PML+G}}_{nn^{\prime}}$ are defined just as in Eqs. (71) and (72), respectively. Now that we have included all the gain components into the loss operators, the gain is completely described by incoherent pumping terms (proportional to $\tilde{\chi}^{\mathrm{PML+G},+}_{nn^{\prime}}$), rather than negative-energy contributions in the effective Hamiltonian. The coupling constants for the interaction with a QE would have the same form as in Eq. (75), but with the accordingly symmetrized modes $\tilde{\bm{E}}^{\mathrm{sym}}_{n}(\bm{r}_{\mathrm{QE}})/\tilde{\bm{B}}^{\mathrm{sym}}_{n}(\bm{r}_{\mathrm{QE}})$. A schematic representation of the quantum description in this simplified setting is shown in Fig. 2(b). In the case of a resonator with a Drude-Lorentz permittivity and non-dispersive permeability, we can also immediately use the formula obtained from the classical QNM theory in Eq. (20) to describe the driving of the modes by the classical coherent field $\bm{E}_{\mathrm{b}}(\bm{r},t)$. In the quantum description, we would introduce an additional driving term $\hat{H}_{\mathrm{drive}}(t)$ in the effective Hamiltonian given by

### v3

@@ III Quantization of the Regularized Quasinormal Modes
$-(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg]\ .$
@@
$\cdot(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg]\ ,$
showing that it is not altered by the presence of the PMLs and reflects the correlation function of the real system without PMLs. In this way, the modal theory reproduces the correct commutators and vacuum correlation functions of the exact theory without the unphysical contribution from the gain components of the PMLs. The magnetic field operator can be expanded using the same bosonic modal operators:
@@ 2. Combination of gain/loss operators
[…]
@@
in which the symmetrized modes are as defined in Eq. (45) with the replacement $P^{\lambda,\mathrm{L}}\rightarrow P^{\lambda}$. From this electric field operator, we would obtain the vacuum correlation function $\langle\hat{E}_{i}(\bm{r})\hat{E}_{j}(\bm{r}^{\prime})\rangle=({\hbar\mu_{0}}/{\pi})\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\,\Im(G_{ij}(\bm{r},\bm{r}^{\prime},\omega))$, and therefore the quantum effects caused by the presence of the gain regions and embodied by the additional term in Eq. (34) are neglected. However, the gain still affects the correlation function as it alters the modes and the Green’s function of the system. Such an approach would not be appropriate to describe systems exhibiting strong gain, as in the context of lasing or exceptional points, where the combined $P$-matrix $P^{\lambda}_{nn^{\prime}}$ may fail to be positive-definite; in this circumstance, the required mode symmetrization cannot be carried out. It should be understood that such behaviour is not a failure exclusive to the modal formalism, but rather a general failure that reflects the underlying approximations of the approach. Indeed, the case in which the $P$-matrix becomes non-positive-definite is one in which the photonic local density of states, proportional to $\Im(G_{ii}(\bm{r},\bm{r},\omega))$, becomes non-positive for some $\bm{r}$ and $\omega$, which marks a fundamental breakdown of the aforementioned approximation of the vacuum correlation function via the imaginary part of the Green’s function, since $\langle\hat{E}_{i}(\bm{r})\hat{E}_{i}(\bm{r})\rangle$ can become negative and this is unphysical.
@@ 3. Combination of all four operators
It is also worth mentioning that in our approach, the limit of lossless/gainless materials (i.e., $\zeta_{\mathrm{i}}\rightarrow 0$) can be taken from the beginning, since we map the problem onto a compact domain $\Omega\subset\mathbb{R}^{3}$. This means that in this case, we can interchange the limit and the integration in the symmetrization matrix elements $P^{\lambda,\mathrm{L}}_{nn^{\prime}}$, and consequently only the PML region $\Omega_{\mathrm{PML}}$ contributes, which corresponds to the radiation loss in Ref. franke2020fluctuation.
@@ IV Effective Lindblad Master Equation
In this section, we derive an effective Lindblad master equation for the electromagnetic modes and the additional coupling to a QE, described as a two-level system (TLS). We will start by considering the dynamics of the modes only in the most general scenario, in which they are characterized by the operators $\hat{a}_{n,\lambda,X}$. In the Heisenberg picture, the dynamics of each modal operator is determined by the Heisenberg equation of motion $\mathrm{d}\hat{O}/\mathrm{d}t=\mathrm{i}[\hat{H},\hat{O}$$\hbar$, leading to franke2022coupled
$\frac{\mathrm{d}\hat{a}_{n,\lambda,\mathrm{L}}}{\mathrm{d}t}$ $=\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\hat{a}_{n,\lambda,\mathrm{L}}\right]-\chi^{\lambda,\mathrm{L},-}_{nn^{\prime}}\hat{a}_{n^{\prime},\lambda,\mathrm{L}}+\hat{F}_{n,\lambda,\mathrm{L}}(t)\ ,$ (60)
@@
$\chi^{\lambda,X,-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}(P^{\lambda,X})^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})P^{\lambda,X}_{n^{\prime\prime}n^{\prime\prime\prime}}(P^{\lambda,X})^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ ,$ (64)
@@
The $\hat{F}_{n,\lambda,X}(t)$ are additional operator terms that still depend on the continuum operators $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$, and which cannot be expressed simply in terms of the modal operators. From the classical QNM theory, it is clear that the amplitudes of the freely evolving modes are oscillating with the complex frequencies $\tilde{\omega}_{n}$, as can be seen in Eq. (16), meaning the expectation values of the modal operators should satisfy $\langle\hat{\xi}_{n,\lambda,\mathrm{L}}(t)\rangle=\langle\hat{\xi}_{\mu,\lambda,\mathrm{L}}(0)\rangle\mathrm{e}^{-\mathrm{i}\tilde{\omega}_{n}t}$ and $\langle\hat{\xi}_{n,\lambda,\mathrm{G}}(t)\rangle=\langle\hat{\xi}_{n,\lambda,\mathrm{G}}(0)\rangle\mathrm{e}^{+\mathrm{i}\tilde{\omega}^{*}_{n}t}$. From this requirement, it is clear that $\langle\hat{F}^{(\dagger)}_{n,\lambda,X}(t)\rangle=0$. Furthermore, the peaked structure was factored out in the spectral representations of the operators $\hat{F}_{n,\lambda,X}(t)$ due to terms of the form $(\tilde{\omega}_{n}-\omega)A_{n}(\omega)=\varepsilon_{0}c^{2}/\tilde{\omega}_{n}$. We thus make a Markov approximation and treat them as white quantum noise forces induced by the dissipation/gain and that ensure the preservation of the commutation relations; that is, they satisfy $\left[\hat{F}_{n,\lambda,\mathrm{X}}(t),\hat{F}^{\dagger}_{n^{\prime},\lambda,X}(t^{\prime})\right]=2\chi^{\lambda,X,-}_{nn^{\prime}}\delta(t-t^{\prime})$. This procedure is similar to the reaction coordinate mapping, in which the coupling of a TLS to a structured bath having a strongly peaked spectral density is mapped to an interaction with a discrete mode and a residual continuum that can then be treated within the Markov approximation binder2018thermodynamics; anto2021capturing. Note that in the equation for the modal loss operators $\hat{a}_{n,\lambda,\mathrm{L}}$, only a single loss-induced noise force $\hat{F}_{m,\lambda,\mathrm{L}}$ appears that counteracts the loss term described by $\chi_{nn^{\prime}}^{\lambda,\mathrm{L},-}$, caused by the overall loss in the system and the gain components of the PMLs, which we have included in the loss operators [see Eq. (III)]. The gain components of the PMLs are thus only treated as a reduction of the loss in the system, as they cannot induce any real physical gain to the modes in terms of pumping. This is different to how the real gain components of the resonator have to be treated if the gain and loss operators are combined, as in this case the gain introduces pumping of the modes franke2022coupled. Where the loss and gain operators are treated separately, the pumping of the modes due to the gain components of the resonator is described by negative-energy contributions in the effective Hamiltonian of Eq. (62), associated to the modal gain operators.
@@
By invoking a quantum analogue of Itô-Stratonovich calculus, following the same procedure as in Refs.gardiner1985input; franke2019quantization; franke2022coupled, an effective Markovian master equation, corresponding to the Heisenberg-Langevin equations in Eqs. (60) and (61), can be obtained in the form
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\rho\right]+\chi^{\lambda,\mathrm{L},-}_{nn^{\prime}}\mathcal{D}^{\lambda,\mathrm{L}}_{nn^{\prime}}\rho+\chi^{\lambda,\mathrm{G},-}_{nn^{\prime}}\mathcal{D}^{\lambda,\mathrm{G}}_{nn^{\prime}}\rho\ ,$ (65)
@@
$\mathcal{D}^{\lambda,X}_{nn^{\prime}}\rho=2\hat{a}_{n^{\prime},\lambda,X}\rho\hat{a}_{n,\lambda,X}^{\dagger}-\{\rho,\hat{a}_{n,\lambda,X}^{\dagger}\hat{a}_{n^{\prime},\lambda,X}\}\ .$ (66)
This is the effective master equation that governs the quantum dynamics of the modes for the most general case, featuring decomposed electric/magnetic modal operators. Note that the loss/gain present in the system necessitates a symmetrization of the modes and consequently induces a coupling between the bosonic modes in both the effective Hamiltonian and the Lindblad dissipators (i.e., coherent and dissipative coupling wang2020dissipativemag). The cross terms in the Lindblad dissipators capture interference effects between different decay channels, as the coupled modes dissipate into the same bath. Such cross terms were also derived for a Jaynes-Cummings model where both the QE and the cavity mode dissipate into the same bath, leading to coupled dissipators between the QE and mode operators to describe Fano-like effects yamaguchi2021theory.
[…]
@@
in which $|\sigma^{\pm}=\ket{\mathrm{e/g}}\bra{\mathrm{g/e}}$ are the usual TLS raising and lowering operators. Note that whilst we treat both possibilities in Eq. (67) for completeness, the distinct selection rules pertaining to electric and magnetic dipole transitions will ensure that only one of $\bm{d}$ or $\bm{m}$ is non-zero in a given physical application. By using the modal expansion of the electric field operator in Eq. (47) and following the same steps as before, the master equation thereby obtained possesses the same form as in Eq. (65), but with an effective Hamiltonian
@@
and we have applied the rotating-wave approximation. It can be appreciated that the gain modal creation operators $\hat{a}^{\dagger}_{n,\lambda,\mathrm{G}}$ couple to $\sigma^{+}$ in the rotating-wave approximation since they appear in the positive-frequency parts of the fields in Eqs. (47) and (51). If the QE is subject to additional non-radiative decay channels or dephasing processes, the corresponding dissipators should be added to the master equation in Eq. (65). Fig. 2(a) offers a schematic representation of the quantum description for an exemplary resonator-QE system, based on the master equation in this general setting, where the action of the operators and all the different couplings and decay channels are labeled. Note that even though the master equation does not contain coupled dissipators between the QE and mode operators, as derived in Ref. yamaguchi2021theory, it is in principle capable of describing coherent Fano interference effects occurring in the QE decay, as the coupling of the QE with the complete set of modes captures its interaction with the electromagnetic environment completely. Such cross terms would appear if a subset of the modes were traced out, as both the QE and the remaining modes couple to these modes.
If the electric and magnetic modal operators are combined, the derivation of the master equation would proceed in a similar manner as before, with the result having the same form as Eq. (65), but with the replacements $\chi^{\lambda,X,\pm}\rightarrow\chi^{X,\pm}$, which are exactly defined as in Eqs. (63) and (64), but with the replacement $P^{\lambda,X}\rightarrow P^{X}$.
@@
$\frac{\mathrm{d}\hat{a}_{n}}{\mathrm{d}t}=\frac{\mathrm{i}}{\hbar}\left[\hat{H}_{\mathrm{eff}},\hat{a}_{n}\right]-\tilde{\chi}_{nn^{\prime}}^{\mathrm{L},-}\hat{a}_{n^{\prime}}+\hat{F}_{n,\mathrm{L}}(t)+\hat{F}^{\dagger}_{n,\mathrm{G}}(t)\ ,$ (70)
@@
$\tilde{\chi}^{X,-}_{nn^{\prime}}$ $=\frac{\mathrm{i}}{2}P^{-\frac{1}{2}}_{nn^{\prime\prime}}(\tilde{\omega}_{n^{\prime\prime}}-\tilde{\omega}_{n^{\prime\prime\prime}}^{*})P^{X}_{n^{\prime\prime}n^{\prime\prime\prime}}P^{-\frac{1}{2}}_{n^{\prime\prime\prime}n^{\prime}}\ ,$ (73)
@@
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{i}}{\hbar}\left[H_{\mathrm{eff}},\rho\right]+\tilde{\chi}_{nn^{\prime}}^{\mathrm{L},-}\mathcal{D}^{\mathrm{L}}_{nn^{\prime}}\rho+\tilde{\chi}_{nn^{\prime}}^{\mathrm{G},-}\tilde{\mathcal{D}}^{\mathrm{G}}_{nn^{\prime}}\rho\ ,$ (74)
in which the dissipators $\mathcal{D}^{\mathrm{L}}_{nn^{\prime}}$ are defined just as in Eqs. (66), and $\tilde{\mathcal{D}}^{\mathrm{G}}_{nn^{\prime}}$ as
$\tilde{\mathcal{D}}^{\mathrm{G}}_{nn^{\prime}}=2\hat{a}_{n}^{\dagger}\rho\hat{a}_{n^{\prime}}-\{\rho,\hat{a}_{n^{\prime}}\hat{a}_{n}^{\dagger}\}\ .$ (75)
In this case, any gain of the resonator is completely described by incoherent pumping terms (proportional to $\tilde{\chi}^{\mathrm{G},+}_{nn^{\prime}}$), rather than negative-energy contributions in the effective Hamiltonian. The coupling constants for the interaction with a QE would have the same form as in Eq. (69), but with the accordingly symmetrized modes $\tilde{\bm{E}}^{\mathrm{sym}}_{n}(\bm{r}_{\mathrm{QE}})/\tilde{\bm{B}}^{\mathrm{sym}}_{n}(\bm{r}_{\mathrm{QE}})$. A schematic representation of the quantum description in this simplified setting is shown in Fig. 2(b). In the case of a resonator with a Drude-Lorentz permittivity and non-dispersive permeability, we can also immediately use the formula obtained from the classical QNM theory in Eq. (16) to describe the driving of the modes by the classical coherent field $\bm{E}_{\mathrm{b}}(\bm{r},t)$. In the quantum description, we would introduce an additional driving term $\hat{H}_{\mathrm{drive}}(t)$ in the effective Hamiltonian given by
@@
In practice, an important issue surrounding the numerical solution of the quantum master equation, Eq. (65) or Eq. (74), is the exponential scaling of the Hilbert space dimension with the number of modes retained to ensure effective completeness, which may quickly render quantum dynamic simulations infeasible unless suitably optimized, numerical linear algebra and parallel computation strategies are leveraged.